# Multi-student Diffusion Distillation for Better One-step Generators

### v1

@@ Abstract
Diffusion models achieve high-quality sample generation at the cost of a lengthy multistep inference procedure. To overcome this, diffusion distillation techniques produce student generators capable of matching or surpassing the teacher in a single step. However, the student model’s inference speed is limited by the size of the teacher architecture, preventing real-time generation for computationally heavy applications. In this work, we introduce Multi-Student Distillation (MSD), a framework to distill a conditional teacher diffusion model into multiple single-step generators. Each student generator is responsible for a subset of the conditioning data, thereby obtaining higher generation quality for the same capacity. MSD trains multiple distilled students, allowing smaller sizes and, therefore, faster inference. Also, MSD offers a lightweight quality boost over single-student distillation with the same architecture. We demonstrate MSD is effective by training multiple same-sized or smaller students on single-step distillation using distribution matching and adversarial distillation techniques. With smaller students, MSD gets competitive results with faster inference for single-step generation. Using $4$ same-sized students, MSD sets a new state-of-the-art for one-step image generation: FID $1.20$ on ImageNet-64$\times$64 and 8.20 on zero-shot COCO2014.  [^1] Project page: https://research.nvidia.com/labs/toronto-ai/MSD/
@@ 1 Introduction
Diffusion models are the dominant generative model in image, audio, video, 3D assets, protein design, and more [20, 28, 6, 2, 47]. They allow different conditioning inputs – such as class labels, text, or images – and achieve high-quality generations. However, their inference process typically requires hundreds of model evaluations – with an often slow and bulky network – for a single sample. This procedure costs millions of dollars per day [65, 15]. It also prohibits applications requiring rapid synthesis, such as augmented reality. Real-time, low-cost, and high-quality generation will have huge financial and operational impacts while enabling new usage paradigms.
There has been a flurry of work on diffusion-distillation techniques to address the slow sampling of diffusion models [41, 64, 75]. Inspired by knowledge distillation [19], these methods use the trained diffusion model as a teacher and optimize a student model to match its generated output in as few as a single step. However, most of these methods use the same student and teacher architecture, preventing real-time generation for applications with bulky networks, such as video synthesis [6]. While reducing model size can reduce latency, it typically yields worse generation quality, thus presenting a speed-to-quality tradeoff dilemma with existing distillation methods.
@@
When distilling into the same-sized students, MSD has the flexibility of conceptually applying to any distillation method for a performance boost. In addition, MSD allows distilling into multiple smaller students for reduced single-step generation time. Using smaller students prevents one from initializing a student from teacher weights, which we solve by adding a relatively lightweight score-matching pretraining stage before distillation (Sec. 4.3).
We validate our approaches by applying MSD to distill the teacher into $4$ same-sized students using distribution matching and adversarial distillation procedures [75, 74]. The resulting students collectively outperform single-student counterparts, setting new state-of-the-art FID scores of $1.20$ on one-step ImageNet-64$\times$64 generation (Tab. 2) and 8.20 on one-step zero-shot COCO2014 generation (Tab. 2). We also distill the same teacher into $4$ smaller students, which achieve a competitive FID of 2.88 on ImageNet (Tab. 2), with $42\%$ less parameters per student.
@@
Figure 2: Samples on high guidance-scale text-to-image generations from the SD v1.5 teacher and different sized students, with full training details in App. B. The same-sized student has comparable quality to the teacher. The smaller student, trained on a subset of dog-related data, achieves faster generation while still having decent qualities. The same-sized student is trained with DM stage only, whereas the smaller student is trained with TSM and DM stages (see Fig. 3). See additional samples in Fig. 11.
@@
  • Demonstrating the effectiveness of MSD by training multiple same sized students using SOTA distillation techniques [75, 74] in Sec. 4.2, resulting in new record FID scores in ImageNet-64$\times$64 (Sec. 5.2) and zero-shot text-to-image generation (Sec. 5.3).
@@ 2 Related Work
The methods above require the student to follow the teacher’s trajectories. Instead, a recent line of works aims to only match the distribution of the student and teacher output via variational score distillation [75, 74, 56, 69, 43, 84, 45]. The adversarial loss [14], often combined with the above techniques, has been used to enhance the distillation performance further [68, 80, 59, 60, 66, 71, 33, 27]. Although MSD is conceptually compatible and offers a performance boost to all of these distillation methods, in this work, we demonstrate two specific techniques: distribution matching [75] and adversarial distillation [74].
Mixture of experts training and distillation. Mixture of Experts (MoE), first proposed in [24], has found success in training large-scale neural networks [61, 30, 13, 31, 7]. Distilling a teacher model into multiple students was explored by Hinton et al. [19], and after that, further developed for supervised learning [10, 46, 9] and language modeling [70, 29, 85]. Although several works [21, 48, 1] have proposed MoE training for generative adversarial networks, they train the MoE from scratch. This requires carefully tuning the multi-expert adversarial losses. A recent work [83] proposes to distill a pretrained diffusion model into an MoE for policy learning, which shares similar motivations with our work. However, to the best of our knowledge, MSD is the first method to distill multi-step teacher diffusion models into multiple one-step students for image generation.
@@ 4.1 Distilling into multiple students
The partition of $\mathcal{Y}$ into $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ determines the input condition groups for each student. As a starting point, we make the following three simplifications for choosing a partition:
@@
  • Equal size: Since students have the same architecture, the partitions $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ should be of equal size that require similar model capacity.
@@
The data filtering function $F$ determines the training subset data $\mathcal{D}_{k}$ from $\mathcal{Y}_{k}$. For example, a vanilla filtering strategy could set $F(\mathcal{D},\mathcal{Y}_{k})=\mathcal{D}_{k}:=\mathcal{D}|\mathcal{Y}_{k}$, where $\mathcal{D}|\mathcal{Y}_{k}$ denotes the subset of the training dataset $\mathcal{D}$ that contains the condition $\mathcal{Y}_{k}$. Empirically, we found that this filtering works in most cases, although sometimes a different approach is justified, as demonstrated in Sec. 4.2.
@@ 4.2 MSD with distribution matching
where we recall that $\bm{\mu}_{\text{teacher}}$ is the teacher diffusion model, $G_{k}^{(i)}$ is the $k$-th student generator at the $i$-th stage, $F$ is the data filtering function, $\mathcal{D}$ is the training data, and $\mathcal{Y}_{k}$ is the set of labels that student $k$ is responsible of. The first stage $\text{Distill}_{\text{DM}}$ uses distribution matching with either a complemented regression loss or the TTUR, with details in Sec. 3.2. These two methods achieve optimal training efficiency among other best-performing single-step distillation methods [69, 84, 27] without an adversarial loss, with a detailed comparison in App. A.3. The second stage $\text{Distill}_{\text{ADM}}$ adds an additional adversarial loss (details in Sec. 3.3), by introducing a small prediction head on the fake score model. This introduces minimal additional computational overhead and allows resuming from the first stage checkpoint.
@@
For the first stage data filtering $F_{\text{DM}}$, we propose $F_{\text{DM}}(\mathcal{D}_{\text{DM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{paired}},\mathcal{C}|\mathcal{Y}_{k})$, where $\mathcal{C}|\mathcal{Y}_{k}$ denotes the subset of condition inputs $\mathcal{C}$ that contains $\mathcal{Y}_{k}$. In other words, we sample all input conditions only on the desired partition for the KL loss but use the whole paired dataset for the regression loss. This special filtering is based on the observation that the size of $\mathcal{D}_{\text{paired}}$ critically affects the terminal performance of DMD distillation: using fewer pairs causes mode collapse, whereas using more pairs challenges the model capacity. Naïvely filtering paired datasets by partition reduces the paired dataset size for each student and leads to worse performance, as in our ablation in App. A.2. Instead of generating more paired data to mitigate this imbalance, we simply reuse the original paired dataset for the regression loss. This is remarkably effective, which we hypothesize is because paired data from other input conditions provides effective gradient updates to the shared weights in the network.
For the second stage, we stick to the simple data filtering $F_{\text{ADM}}(\mathcal{D}_{\text{ADM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{real}}|\mathcal{Y}_{k},\mathcal{C}|\mathcal{Y}_{k})$, so that both adversarial and KL losses focus on the corresponding partition, given that each student has enough mode coverage from the first stage.
@@ 4.3 Distilling smaller students from scratch
Via the frameworks presented in the last two sections, MSD enables a performance upgrade over alternatives for one student with the same model architecture. In this section, we investigate training multiple students with smaller architectures – and thus faster inference time – without impairing much performance. However, this requires distilling into a student with a different architecture, preventing initialization from pretrained teacher weights. Training a single-step student from scratch has previously been difficult [69], and we could not obtain competitive results with the simple pipeline in Eq. 7. Therefore, we propose an additional pretraining phase $\text{Distill}_{\text{TSM}}$, with TSM denoting Teacher Score Matching, to find a good initialization for single-step distillation. TSM employs the following score-matching loss:
@@ 5 Experiments
To focus on the performance boost from multi-student distillation, we applied minimal changes to the hyperparameters used by Yin et al. [75, 74] for their distribution matching distillation implementations. More details on training and evaluation can be found in the App. B and C.
@@ 5.1 Toy Experiments
In Fig. 4, we show the sample density of MSD with DM stage for a 2D toy experiment, where the real data distribution has $8$ classes, and each class is a mixture of $8$ Gaussians. We used a simple MLP with EDM schedules to train the teacher and then distill into $1$, $2$, $4$, and $8$ students for comparison. From the displayed samples and the $\ell_{1}$ distance from teacher generation, we observe that the collective generation quality increases as the number of students increases.
@@ 5.2 Class-conditional Image Generation
Student architecture the same as the teacher: We trained $K=4$ students using the MSD framework and the EDM [25] teacher on class-conditional ImageNet-64$\times$64 generation. We applied the simplest strategy for splitting classes among students: Each student is responsible for $250$ consecutive classes in numerical order. We compare the performance with previous methods and display the results in Tab. 2. Our DM stage, which uses the complementary regression loss, surpasses the one-student counterpart DMD [75], achieving a modest drop of $0.25$ in FID score, making it a strong competitor in single-step distillation without an adversarial loss. We then took the best pretrained checkpoints and trained with the ADM stage. The resulting model achieved the current state-of-the-art FID score of $1.20$. It surpasses even the EDM teacher, StyleGAN-XL [57], the multi-step RIN [23] due to the adversarial loss. Fig. 5(a) and (b) display a comparison of sample generations, showing that our best students have comparable generation quality as the teacher.
Table 1: Comparing class-conditional generators on ImageNet-64$\times$64. We measure generation speed with the number of function evaluations (NFE) and quality with FID.
@@
Student architecture smaller than the teacher: Next, we trained $4$ smaller student models with the prepended teacher score matching (TSM) stage from Sec. 4.3. This achieved a $42\%$ reduction in model size and a $7\%$ reduction in latency, with a slight degradation in FID score, offering a flexible framework to increase generation speed by reducing student size, and increasing generation quality by training more students. Fig. 5(c) displays sample generation from these smaller students, whereas Fig. 5(d) shows sample generations from an even smaller set of students, with a $71\%$ percent reduction in model size and a $23\%$ percent reduction in latency. We observed slightly degraded but still competitive generation qualities. Using more and larger students will further boost performance, as shown by ablations in Sec. 5.4 and App. A.4. Smaller students without the TSM stage fail to reach competitive performance, indicating its necessity.
@@
Figure 5: Sample generations on ImageNet-64$\times$64 from the teacher and different sized students, with architecture and latency details in App. B. The same-size students have comparable or better generation quality than the teacher. Smaller students achieve faster generation while still having decent qualities. Same-sized students are trained with DM and ADM stages, whereas smaller students are trained with all three stages (see Fig. 3). See additional samples in Fig. 9 and Fig. 10.
@@ 5.3 Text-to-Image Generation
Student architecture the same as the teacher: We evaluated the performance of text-to-image generation using the MS-COCO2014 [34] evaluation dataset. We distilled $4$ students from Stable Diffusion (SD) v1.5 [54] on a 5M-image subset of the COYO dataset [8]. For splitting prompts among students, we again employed a minimalist design: pass the prompts through the pre-trained SD v1.5 text encoder, pool the embeddings over the temporal dimension, and divide into $4$ subsets along $4$ quadrants. We trained with a classifier-free guidance (CFG) scale of $1.75$ for best FID performance. Tab. 2 compares the evaluation results with previous methods. Our baseline method with only the DM stage again achieved a performance boost with a $0.48$ drop in FID over the single-student counterpart DMD2 without adversarial loss [74]. Continuing the ADM stage from the best checkpoint yielded a terminal FID of 8.20, again surpassing the single-student counterpart and achieving the current state-of-the-art FID score. In addition, for better visual quality, we also train with a larger CFG scale of $8$, and display corresponding samples in Fig. 2(b) and App. D.2.
Student architecture smaller than the teacher: As a preliminary exploration, with the prepended teacher score matching (TSM) stage, we train a $83\%$ smaller and $5\%$ faster student on a dog-related prompt subset of COYO (containing 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts). We trained with a CFG scale of $8$ and display the samples in Fig. 2. We observed fair generation quality despite a significant drop in model size. Improved training is likely to obtain better sample quality and generalization power.

### v2

@@ Abstract
Diffusion models achieve high-quality sample generation at the cost of a lengthy multistep inference procedure. To overcome this, diffusion distillation techniques produce student generators capable of matching or surpassing the teacher in a single step. However, the student model’s inference speed is limited by the size of the teacher architecture, preventing real-time generation for computationally heavy applications. In this work, we introduce Multi-Student Distillation (MSD), a framework to distill a conditional teacher diffusion model into multiple single-step generators. Each student generator is responsible for a subset of the conditioning data, thereby obtaining higher generation quality for the same capacity. MSD trains multiple distilled students, allowing smaller sizes and, therefore, faster inference. Also, MSD offers a lightweight quality boost over single-student distillation with the same architecture. We demonstrate MSD is effective by training multiple same-sized or smaller students on single-step distillation using distribution matching and adversarial distillation techniques. With smaller students, MSD gets competitive results with faster inference for single-step generation. Using $4$ same-sized students, MSD significantly outperforms single-student baseline counterparts and achieves remarkable FID scores for one-step image generation: $1.20$ on ImageNet-64$\times$64 and $8.20$ on zero-shot COCO2014.  [^1] Project page: https://research.nvidia.com/labs/toronto-ai/MSD/
@@ 1 Introduction
Diffusion models are the dominant generative model in image, audio, video, 3D assets, protein design, and more [20, 28, 6, 2, 47]. They allow conditioning inputs – such as class labels, text, or images – and achieve high-quality generated outputs. However, their inference process typically requires hundreds of model evaluations – with an often slow and bulky network – for a single sample. This procedure costs millions of dollars per day [66, 15] and prohibits applications requiring rapid synthesis, such as augmented reality. Real-time, low-cost, and high-quality generation will have enormous financial and operational impacts while enabling new usage paradigms.
There has been a flurry of work on diffusion-distillation techniques to address diffusion model’s slow sampling [41, 65, 76]. Inspired by knowledge distillation [19], these methods use the trained diffusion model as a teacher and optimize a student model to match its output in as little as a single step. However, most diffusion distillation methods use the same student and teacher architecture. This prevents real-time generation for applications with bulky networks, such as video synthesis [6]. While reducing model size can reduce inference time, it typically yields worse generation quality, thus presenting a speed-to-quality tradeoff dilemma with existing distillation methods.
@@
When distilling into the same-sized students, MSD has the flexibility of conceptually applying to any distillation method for a performance boost. In addition, MSD allows distilling into multiple smaller students for reduced single-step generation time. Using smaller students prevents one from initializing a student from teacher weights, posing an additional technical challenge. We solve this challenge by adding a relatively lightweight score-matching pretraining stage before distillation (Sec. 4.3) and demonstrating its necessity and efficiency via extensive experiments.
We validate our approaches by applying MSD to distill the teacher into $4$ same-sized students using distribution matching and adversarial distillation procedures [76, 75]. The resulting students collectively outperform single-student counterparts, achieving close to state-of-the-art FID scores of $1.20$ on one-step ImageNet-64$\times$64 generation (Tab. 2) and $8.20$ on one-step zero-shot COCO2014 generation (Tab. 2). We also distill the same teacher into $4$ smaller students, which achieve a competitive FID of $2.88$ on ImageNet (Tab. 2), with $42\%$ less parameters per student.
@@
Figure 2: Samples on high guidance-scale text-to-image generations from the SD v1.5 teacher and different sized students, with full training details in App. D. The same-sized student has comparable quality to the teacher. The smaller student, trained on a subset of dog-related data, achieves faster generation while still having decent qualities. The same-sized student is trained with DM stage only, whereas the smaller student is trained with TSM and DM stages (see Fig. 3). See additional samples in Fig. 12.
@@
  • Demonstrating the effectiveness of MSD by training multiple same sized students using SOTA distillation techniques [76, 75] in Sec. 4.2, surpassing single-student counterparts in FID scores for ImageNet-64$\times$64 (Sec. 5.2) and zero-shot text-to-image generation (Sec. 5.3).
@@ 2 Related Work
The methods above require the student to follow the teacher’s trajectories. Instead, a recent line of works aims to only match the distribution of the student and teacher output via variational score distillation [76, 75, 56, 70, 43, 88, 45, 86]. The adversarial loss [14], often combined with the above techniques, has been used to enhance the distillation performance further [69, 82, 59, 60, 67, 72, 33, 27, 87]. Although MSD is conceptually compatible and offers a performance boost to all of these distillation methods, in this work, we demonstrate two specific techniques: distribution matching [76] and adversarial distillation [75].
Mixture of experts training and distillation. Mixture of Experts (MoE), first proposed in [24], has found success in training very large-scale neural networks [62, 30, 13, 31, 7]. Distilling a teacher model into multiple students was explored by Hinton et al. [19], and after that, has been further developed for supervised learning [10, 46, 9] and language modeling [71, 29, 89]. Although several works [21, 48, 1] have proposed MoE training schemes for generative adversarial networks, they train the MoE from scratch. This requires carefully tuning the multi-expert adversarial losses. eDiff-I [3] uses different experts in different denoising timesteps for a multi-step diffusion model. A recent work [85] proposes to distill a pretrained diffusion model into an MoE for policy learning, which shares similar motivations with our work. However, to the best of our knowledge, MSD is the first method to distill multi-step teacher diffusion models into multiple one-step students for image generation.
@@ 4.1 Distilling into multiple students
The partition of $\mathcal{Y}$ into $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ determines the input condition groups each student is responsible for. As a starting point, we make the following three simplifications for choosing a partition:
@@
  • Equal size: Since students have the same architecture, the partitions $\{\mathcal{Y}_{k}\}_{k=1}^{K}$ should be of equal size and require similar model capacity.
@@
The data filtering function $F$ determines the training subset data $\mathcal{D}_{k}$ from $\mathcal{Y}_{k}$. For example, a vanilla filtering strategy could set $F(\mathcal{D},\mathcal{Y}_{k})=\mathcal{D}_{k}:=\mathcal{D}_{\mathcal{Y}_{k}}$, where $\mathcal{D}_{\mathcal{Y}_{k}}$ denotes the training data subset $\mathcal{D}$ containing the desired condition $\mathcal{Y}_{k}$. Empirically, we found that this filtering works in most cases, although sometimes a different approach is justified, as demonstrated in Sec. 4.2.
@@ 4.2 MSD with distribution matching
where we recall that $\bm{\mu}_{\text{teacher}}$ is the teacher diffusion model, $G_{k}^{(i)}$ is the $k$-th student generator at the $i$-th stage, $F$ is the data filtering function, $\mathcal{D}$ is the training data, and $\mathcal{Y}_{k}$ is the set of labels that student $k$ is responsible of. The first stage $\text{Distill}_{\text{DM}}$ uses distribution matching with either a complemented regression loss or the TTUR, with details in Sec. 3.2. These two methods achieve optimal training efficiency among other best-performing single-step distillation methods [70, 88, 27] without an adversarial loss, with a detailed comparison in App. B.3. The second stage $\text{Distill}_{\text{ADM}}$ adds an additional adversarial loss (detailed in Sec. 3.3), which introduces minimal additional computational overhead and allows resuming from the first stage checkpoint.
@@
For the first stage data filtering $F_{\text{DM}}$, we propose $F_{\text{DM}}(\mathcal{D}_{\text{DM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{paired}},\mathcal{C}_{\mathcal{Y}_{k}})$, where $\mathcal{C}_{\mathcal{Y}_{k}}$ denotes the subset of condition inputs $\mathcal{C}$ that contains $\mathcal{Y}_{k}$. In other words, we sample all input conditions only on the desired partition for the KL loss but use the whole paired dataset for the regression loss. This special filtering is based on the observation that the size of $\mathcal{D}_{\text{paired}}$ critically affects the terminal performance of DMD distillation: using fewer pairs causes mode collapse, whereas using more pairs challenges the model capacity. Naïvely filtering paired datasets by partition reduces the paired dataset size for each student and leads to worse performance, as in our ablation in App. B.2. Instead of generating more paired data to mitigate this imbalance, we simply reuse the original paired dataset for the regression loss. This is remarkably effective, which we hypothesize is because paired data from other input conditions provides effective gradient updates to the shared weights in the network.
For the second stage, we stick to the simple data filtering $F_{\text{ADM}}(\mathcal{D}_{\text{ADM}},\mathcal{Y}_{k})=(\mathcal{D}_{\text{real},\mathcal{Y}_{k}},\mathcal{C}_{\mathcal{Y}_{k}})$, so that both adversarial and KL losses focus on the corresponding partition, given that each student has enough mode coverage from the first stage.
@@ 4.3 Distilling smaller students from scratch
Via the frameworks presented in the last two sections, MSD enables a performance upgrade over alternatives for one student with the same model architecture. In this section, we investigate training multiple students with smaller architectures – and thus faster inference time – without impairing much performance. However, this requires distilling into a student with a different architecture, preventing initialization from pretrained teacher weights. Distilling a single-step student from scratch has previously been difficult [70], and we could not obtain competitive results with the simple pipeline in Eq. 7. Therefore, we propose an additional pretraining phase $\text{Distill}_{\text{TSM}}$, with TSM denoting Teacher Score Matching, to find a good initialization for single-step distillation. TSM employs the following score-matching loss:
@@ 5 Experiments
To focus on the performance boost from multi-student distillation, we applied minimal changes to the hyperparameters used by Yin et al. [76, 75] for their distribution matching distillation implementations. More details on training and evaluation can be found in the App. D and E.
@@ 5.1 Toy Experiments
In Fig. 4, we show the sample density of MSD with DM stage for a 2D toy experiment, where the real data distribution has $8$ classes, and each class is a mixture of $8$ Gaussians. We used a simple MLP with EDM schedules to train the teacher and distill into $1$, $2$, $4$, and $8$ students for comparison. The displayed samples and the $\ell_{1}$ distance from teacher generation show the collective generation quality increases as the number of students increases. App. A.2 additionally shows the results of Consistency Distillation [65] in this setting, demonstrating our framework’s generality.
@@ 5.2 Class-conditional Image Generation
Student architecture the same as the teacher: We trained $K=4$ students using the MSD framework and the EDM [25] teacher on class-conditional ImageNet-64$\times$64 generation. We applied the simplest strategy for splitting classes among students: Each student is responsible for $250$ consecutive classes in numerical order (i.e., $\nicefrac{{1}}{{K}}$ of the $1000$ classes). We compare performance with prior methods and show results in Tab. 2. Our DM stage, which uses the complementary regression loss, surpasses the one-student counterpart DMD [76], achieving a modest $0.25$ drop in FID score, making it a strong competitor in single-step distillation without an adversarial loss. We then took the best pretrained checkpoints and trained with the ADM stage, again outperforming the single-student counterpart DMD2 [75] with an FID score of $1.20$. It surpasses even the EDM teacher, StyleGAN-XL [57], the multi-step RIN [23] due to the adversarial loss. Fig. 5(a) and (b) compare sample generations, showing our best students have comparable generation quality as the teacher.
Table 1: Comparing class-conditional generators on ImageNet-64$\times$64. The number of function evaluations (NFE) for MSD is $1$ as a single student is used at inference for the given input.
@@
Student architecture smaller than the teacher: Next, we trained $4$ smaller student models with the prepended teacher score matching (TSM) stage from Sec. 4.3. This achieved a $42\%$ reduction in model size and a $7\%$ reduction in latency, with a slight degradation in FID score, offering a flexible framework to increase generation speed by reducing student size and increasing generation quality by training more students. Fig. 5(c) displays sample generation from these smaller students, whereas Fig. 5(d) shows sample generations from an even smaller set of students, with a $71\%$ percent reduction in model size and a $23\%$ percent reduction in latency. We observed slightly degraded but still competitive generation qualities. Using more and larger students will further boost performance, as shown by ablations in Sec. 5.4 and App. B.4. Smaller students without the TSM stage fail to reach even proper convergence. Moreover, instead of the TSM stage, we performed post-output distillation on the best single-step checkpoints and observed a significant drop in performance. Hence, the TSM stage is both necessary and efficient.
@@
Figure 5: Sample generations on ImageNet-64$\times$64 from the teacher and different sized students, with architecture and latency details in App. D. The same-size students have comparable or slightly better generation quality than the teacher. Smaller students achieve faster generation while still having decent qualities. Same-sized students are trained with DM and ADM stages, whereas smaller students are trained with all three stages as shown in Fig. 3.
@@ 5.3 Text-to-Image Generation
Student architecture the same as the teacher: We evaluated the performance of text-to-image generation using the MS-COCO2014 [34] evaluation dataset. We distilled $4$ students from Stable Diffusion (SD) v1.5 [54] on a 5M-image subset of the COYO dataset [8]. For splitting prompts among students, we again employed a minimalist design: pass the prompts through the pre-trained SD v1.5 text encoder, pool the embeddings over the temporal dimension, and divide into $4$ disjoint subsets along $4$ quadrants. We trained with a classifier-free guidance (CFG) scale of $1.75$ for best FID performance. Tab. 2 compares the evaluation results with previous methods. Our baseline method with only the DM stage again achieved a performance boost with a $0.48$ drop in FID over the single-student counterpart DMD2 without adversarial loss [75]. Continuing the ADM stage from the best checkpoint yields a terminal FID of $8.20$, surpassing the single-student counterpart and achieving a close-to state-of-the-art FID score. In addition, for better visual quality, we also train with a larger CFG scale of $8$ with corresponding samples in Fig. 2(b) and App. F.2.
Student architecture smaller than the teacher: As a preliminary exploration, with the prepended teacher score matching (TSM) stage, we train a $83\%$ smaller and $5\%$ faster student on a dog-related prompt subset of COYO (containing 1 210 000similar-toabsent1210000$\sim$\$1\,210\,000\$∼ 1 210 000 prompts). We trained with a CFG scale of $8$ and display the samples in Fig. 2. We observed fair generation quality despite a significant drop in model size. Improved training is likely to obtain better sample quality and generalization power. Due to limited computational resources and the complete coverage of the prompt set by the 4-student model, we did not train the full set of students at this size.